# Pairing mechanism and superconductivity in pressurized La5Ni3O11

### v1

@@ Abstract
The discovery of superconductivity (SC) with critical temperature $T_{c}$ above the boiling point of liquid nitrogen in pressurized La3Ni2O7 has sparked a surge of exploration of high-$T_{c}$ superconductors in the Ruddlesden-Popper (RP) phase nickelates. More recently, the RP phase nicklate La5Ni3O11, which hosts layered structure with alternating bilayer and single-layer NiO2 planes, is reported to accommodate SC under pressure, exhibiting a dome-shaped pressure dependence with highest $T_{c}\approx 64$ K, capturing a lot of interests. Here, using density functional theory (DFT) and random phase approximation (RPA) calculations, we systematically study the electronic properties and superconducting mechanism of this material. Our DFT calculations yield a band structure including two nearly decoupled sets of sub- band structures, with one set originating from the bilayer subsystem and the other from the single-layer one. RPA-based analysis demonstrates that SC in this material occurs primarily within the bilayer subsystem exhibiting an $s^{\pm}$ wave pairing symmetry similar to that observed in pressurized La3Ni2O7, while the single-layer subsystem mainly serves as a bridge facilitating the inter-bilayer phase coherence through the interlayer Josephson coupling (IJC). Since the IJC thus attained is extremely weak, it experiences a prominent enhancement under pressure, leading to the increase of the bulk $T_{c}$ with pressure initially. When the pressure is high enough, the $T_{c}$ gradually decreases due to the reduced density of states on the $\gamma$-pocket. In this way, the dome-shaped pressure dependence of $T_{c}$ observed experimentally is naturally understood.
@@
The discovery of superconductivity (SC) with critical temperature $T_{c}\approx 80$ K in the Ruddlesden-Popper (RP) bilayer nickelate La3Ni2O7 under pressure bilayernature has sparked significant experimental and theoretical interest in the field bilayermodel ; huomodulation ; zhang2023electronic ; lechermann2023electronic ; luo2024high ; wu2024superexchange ; cpl_41_7_077402 ; shilenko2023correlated ; yang2023interlayer ; zhang2023trends ; christiansson2023correlated ; shen2023effective ; oh2023type ; liu2023s ; liao2023electron ; yang2023possible ; PhysRevB.110.235155 ; kaneko2024pair ; ouyang2024absence ; heier2024competing ; zhang2024structural ; zhang2024electronic ; tian2024correlation ; ryee2024quenched ; zhang2024strong ; ni_spin_2025 ; lu2024interlayer ; qu2024bilayer ; yang2024strong ; fan2024superconductivity ; sakakibara2024possible ; cao2024flat ; jiang2024pressure ; chen2025charge ; yang_orbital-dependent_2024 ; NPzhang ; Hou_2023 ; PhysRevB.110.134520 ; wang_pressure-induced_2024 ; liu_electronic_2024 ; PhysRevB.110.L180501 ; PhysRevLett.134.076001 ; shao2024possiblehightemperaturesuperconductivitydriven . Subsequently, researchers observed SC in the trilayer nickelate La4Ni3O10 under pressure, and recently detected SC in ambient-pressure (AP) La3Ni2O7 thin films zhu_superconductivity_2024 ; zhang2025superconductivity ; ko_signatures_2025 ; zhou_ambient-pressure_2025 . These findings further underscore the potential of RP-series nickelates as high-$T_{c}$ superconductors PhysRevB.110.014503 ; PhysRevB.111.075140 ; PhysRevB.109.144511 ; li_signature_2024 ; hu2025electronicstructuresmultiorbitalmodels ; experimental . In RP nickelates with the general formula Rn+1NinO3n+1 (R = rare earth), each Ni forms NiO6 octahedron with corner-sharing O atoms. At AP, bilayer La3Ni2O7 and trilayer La4Ni3O10 manifest octahedral tilting and in-plane lattice anisotropy. Upon increasing pressure, both the octahedral tilting and the in-plane lattice constant disparity gradually diminish and eventually crystal structures arrive at higher symmetry ($I4/mmm$ space group) li_identification_2025 .
@@
Furthermore, hybrid RP phase nickelates, formed by alternately stacking different RP phases along the c-axis, also attract widespread investigations, especially hybridization of single-layer La2NiO4, bilayer La3Ni2O7 and trilayer La4Ni3O10 zhang2024magneticcorrelationspairingtendencies ; chen_jpcm ; PhysRevLett.133.146002 ; PhysRevMaterials.8.053401 ; ouyang2025phasediagramskeyfactors . Until now, successful synthesis of 1313 phase (La3Ni2O7)  chen_jpcm ; PhysRevLett.133.146002 and 1212 phase (La5Ni3O11PhysRevMaterials.8.053401 has been achieved. A previous study reported a superconducting transition onset temperature up to 80 K in alternately stacking single-layer La2NiO4 and trilayer La4Ni3O10, namely 1313 phase PhysRevLett.133.146002 . However, it remains unclear whether the observed high-$T_{c}$ comes from the hybrid 1313 phase or $2$ bilayer phase.
@@
Recently, pressure-induced SC with maximal $T_{c}=$ 64 K was observed in hybrid 1212 phase shi2025superconductivityhybridruddlesdenpopperla5ni3o11 . Transport and magnetic torque measurements also revealed a density-wave (DW) transition at ambient and low pressures. In contrast to the previously reported pressure-induced SC in La3Ni2O7 and La4Ni3O10, the DW observed here is notably robust against pressure. Intriguingly, the superconducting phase emerges at around 11.7 GPa and exhibits a dome-like pressure-dependent behavior, in contrast to the previously reported right-triangle-like behavior observed in pressurized li_identification_2025 . Previous random-phase approximation (RPA) based study proposed that $d$-wave is the leading pairing symmetry that originates from the La2NiO4 single-layer in the material zhang2025electronicstructuremagneticpairing . However, recent studies using combined density functional theory (DFT) and dynamic-mean-field-theory (DMFT) provide another possibility that the La2NiO4 single-layer in La5Ni3O11 is nearly Mott-insulating which does not carry SC ouyang2025phasediagramskeyfactors ; labollita2025correlated , consistent with previous insight that the pure La2NiO4 is an antiferromagnetic (AFM) Mott insulator PhysRevB.75.012414 ; PhysRevB.80.144523 . Therefore, a more systematic investigation into the electronic and superconducting properties of the hybrid 1212 phase La5Ni3O11 is urgently needed.
@@
In this paper, we adopt the first-principle DFT calculations to study the electronic structure of La5Ni3O11, followed by a RPA based study to clarify the superconducting mechanism in this material. Based on our DFT band structure, we construct a trilayer $(d_{z^{2}},d_{x^{2}-y^{2}})$- orbital tight-binding (TB) model, in which the hopping integrals between the bilayer subsystem and the single-layer one is very weak, reflecting that the two subsystems are almost decoupled, allowing for separate treatments of the two parts. Our RPA based analysis suggests that the superconducting pairing mainly occurs in the bilayer subsystem, which takes the s±-pattern similar with that in pressurized La3Ni2O7, while the sinlgle-layer subsystem mainly serves as a bridge connecting adjacent superconducting bilayers to establish phase coherence along the c-axis necessary for the bulk SC, through the interlayer Josephson coupling (IJC). As the IJC is extremely weak, it will experience prominent enhancement with the increase of the pressure, leading to enhancement of the $T_{c}$ in the low pressure regime. When the pressure is strong enough, the $T_{c}$ decreases with further enhancement of the pressure due to a reduction in the density of states (DOS) on the $\gamma$-pocket crucial for pairing. Therefore, our study provides a natural understanding of the dome-shaped pressure dependence of the $T_{c}$ observed in the experiment.
@@ Results
Figure 1: (color online) Crystal structure. Side view of RP crystal structures of single-layer (La2NiO4), bilayer (La3Ni2O7) and SL-BL (La5Ni3O11). The blue, grey, and red balls represent lanthanum, nickel, and oxygen atoms.
La5Ni3O11 is a single-layer-bilayer (SL-BL) stacked nickelate, as shown in Fig. 1. La2NiO4 and La3Ni2O7 stack alternatively in RP phase, forming in $Cmmm$ space group at AP and transitioning to $P4/mmm$ space group under high pressure (HP) (shi2025superconductivityhybridruddlesdenpopperla5ni3o11, ). Unlike bulk La3Ni2O7, the out-of-plane Ni-O-Ni angle in La5Ni3O11 remains 180 from AP to HP, leading to the absence of octahedral distortion.
Band structure of SL-BL stacking La5Ni3O11 at 12 GPa is shown in Fig. 2(b), consistent with previous theoretical results ouyang2025phasediagramskeyfactors ; labollita2025correlated . As in most RP phase nickelates, bands around the Fermi level can be described by the Ni-$e_{g}$ sector. The Ni-$d_{z^{2}}$ band of La2NiO4 lies between the bonding and antibonding bands of La3Ni2O7. Compared to band structure at AP (see Fig.[S2]), the bonding state of La3Ni2O7 is metalized by pressure, leaving the $\gamma$ pocket at the Fermi surface (FS), as shown in Fig. 2(d). The emergence of $\gamma$ pocket implies that the SC in La5Ni3O11 may be attributed to La3Ni2O7, since the appearance of SC in La3Ni2O7 is also accompanied by the emergence of the $\gamma$ pocket.
Figure 2: (color online) Band structure of the DFT and six-orbital TB model for La5Ni3O11 at 12 GPa. (a) Schematic of La5Ni3O11 lattice of six-orbital TB model. The dashed line denotes the $(\frac{1}{2},\frac{1}{2})$ translation of RP stacking between bilayer and single-layer. (b) DFT band structure and projected DOS of La5Ni3O11. (c) The TB band structure corresponding to (a). (d) FS in the brillouin zone, with the five pockets labeled. The color scheme in the left half of (d) represents the relative contributions of the $d_{z^{2}}$ and $d_{x^{2}-y^{2}}$ orbital, while the colors scheme on the right half of (d) indicates the relative contributions from Ni atoms in the single-layer and bilayer.
@@
Here $\bm{r}_{i}$ represents the coordinates of site $i$, $\Delta\bm{r}_{x}(\Delta\bm{r}_{y})\in(-2,2)$ represents the hoppings up to the third-nearest neighbor. In this context, $\bm{r}_{i}$ refers to the unit cell index, while a finite $\Delta\bm{r}$ represents the interunit cell separation. The indices $\mu,\nu=1,\cdots,6$ containing the $d_{z^{2}}$ and $d_{x^{2}-y^{2}}$ orbitals of the upper and lower layers of the bilayer subsystem, as well as those of the single-layer subsystem. The SL-BL lattice and corresponding hoppings are depicted in Fig. 2(a), with the hopping integrals $t_{ij}^{\mu\nu}$ are provided in Tab. 1. Note that due to the RP stacking in La5Ni3O11, a $(\frac{1}{2},\frac{1}{2})$ translation exists between the bilayer and single-layer lattice. Fig. 2(c) displays the band structure for TB model, which is in good agreement with DFT results. The obtained FS for TB model reveals five pockets labeled as $\alpha$, $\alpha^{\prime}$, $\beta$, $\gamma$, $\gamma^{\prime}$, as shown in Fig. 2(d). The dominant component orbital of $\gamma$ and $\gamma^{\prime}$ pockets is $d_{z^{2}}$, while the $\alpha$, $\alpha^{\prime}$ and $\beta$ pockets are the mix of $d_{z^{2}}$ and $d_{x^{2}-y^{2}}$ orbitals. From the right half of Fig. 2(d), we can see that the $\alpha^{\prime}$ and $\gamma^{\prime}$ pocket are contributed from single-layer while the other three pockets from bilayer.
Table 1: Hopping parameters and site energies of TB model for La5Ni3O11 at 12 GPa in unit of eV. Here, $x$, $z$ and $xz$ denotes hopping within and between $d_{x^{2}-y^{2}}$, $d_{z^{2}}$ orbitals, respectively. The abbreviations BI, BO, and SI represent intra-bilayer, inter-bilayer, and single-layer in-plane hoppings, respectively, while BS denotes hoppings between bilayer and single-layer units. The TB model includes hopping processes up to the third-nearest neighbor, i.e., [2,0]. $\epsilon$ denotes the on-site energy.
@@
Figure 3: (color online) DFT calculated band and FS characteristics of the bilayer subsystem and single-layer subsystem in La5Ni3O11 (SL-BL system) at 12 GPa. (a) Band structure of the SL-BL system (black lines) and the isolated bilayer (red lines). (b) FS of the bilayer subsystem, showing that the $\alpha$ pocket is primarily composed of $d_{x^{2}-y^{2}}$ orbitals, the $\beta$ pocket contains contributions from both $d_{x^{2}-y^{2}}$ and $d_{z^{2}}$ orbitals, and the $\gamma$ pocket originates mainly from $d_{z^{2}}$ orbitals. (c) Band structure of the SL-BL system (black lines) and the isolated single-layer (red lines). (d) FS of the single-layer subsystem, showing that the $\alpha^{\prime}$ pocket is primarily composed of $d_{x^{2}-y^{2}}$ orbitals, the $\gamma^{\prime}$ pocket originates mainly from $d_{z^{2}}$ orbitals.
@@
As shown in Tab. 1, the coupling between the single-layer and the bilayer is extremely weak, with the strongest hopping integral amplitude between the two parts to be $\left|t_{[0,0]}^{z,BS}\right|=0.0103$ eV. The sole effect of such a weak coupling is to equilibrate the chemical potentials of the two subsystems and thereby adjust the electron filling. Consequently, it is reasonable to treat the two subsystems separately.
@@
The decoupling of the two subsystems at the TB Hamiltonian level is achieved by modifying the summation over $\mu$ and $\nu$ in Eq. 1. Specifically, restricting $\mu,\nu\in(1,\cdots,4)$ and electron filling $n\approx 3.1$ which got from SL-BL system yields the Hamiltonian of the bilayer subsystem, while restricting $\mu,\nu\in(5,6)$ and electron filling $n\approx 1.9$ yields that of the single-layer subsystem. A comparison between the band structure of the bilayer subsystem and that of La5Ni3O11 (SL-BL system) is shown in Fig. 3(a), where it can be seen that the bilayer subsystem reproduces the key features of the bilayer-derived bands in the full system. The three FS pockets of the bilayer, labeled $\alpha$, $\beta$, and $\gamma$, are shown in Fig. 3(b), and are clearly separated from the five pockets observed in Fig. 2(d). Fig. 3(c) shows that the band structure of the single-layer subsystem matches well with the single-layer-derived bands of the full SL-BL system. The two Fermi pockets of the single-layer subsystem, corresponding to the $\alpha^{\prime}$ and $\gamma^{\prime}$ pockets of the full system, are shown in Fig. 3(d).
@@
Figure 4: (color online) (a) FS of the bilayer subsystem at 12 GPa. (b) Distribution of the largest eigenvalue of the spin susceptibility matrix $\chi^{s}(q)$ in the bilayer subsystem for $U=1$ eV and $J_{H}=U/6$. The susceptibility peaks at two inequivalent momenta, denoted as Q1 and Q2, respectively. As shown in (a), Q1 corresponds to a nesting vector between the $\beta$ and $\gamma$ pockets, while Q2 corresponds to a nesting vector between the $\alpha$ and $\beta$ pockets. (c) The largest pairing eigenvalue $\lambda$ of the various pairing symmetries as function of the interaction strength $U$ with fixed $J_{H}=U/6$. (d) The distributions of the leading $s^{\pm}$-wave pairing gap functions on the FS for $U=1.1$ eV.
@@
The separation of the single-layer and bilayer subsystems within the TB framework points toward an underlying layer-selective character in this hybrid RP-phase nickelate. At ambient and low pressures, the incorporation of local dynamical correlations via the DMFT modifies the DFT band structure by removing the FS pockets contributed by the single‐layer subsystem, leaving only the three FS pockets contributed by the bilayer subsystem ouyang2025phasediagramskeyfactors ; labollita2025correlated . This observation indicates that the single-layer subsystem is close to a Mott‐insulating state, with possible AFM order present ouyang2025phasediagramskeyfactors , analogous to the pure La2NiO4 PhysRevB.75.012414 ; PhysRevB.80.144523 . At HP, the Mott gap becomes narrower, permitting a small density of mobile carriers that may suppress the AFM order ouyang2025phasediagramskeyfactors . Indeed, the long‐range magnetic order has not been observed experimentally under HP so far PhysRevMaterials.8.053401 . Regardless of whether residual magnetic order persists or not, the single-layer subsystem remains a “bad metal” and can hardly carry SC ouyang2025phasediagramskeyfactors . Consequently, SC must arise predominantly from the bilayer subsystem. Since DFT+DMFT indicates that electronic correlation in the bilayer subsystem is comparatively weaker, we will focus our subsequent RPA analysis on the bilayer subsystem to study its pairing instability.
@@
We adopt the following multi-orbital Hubbard interaction to investigate the SC drived by electron interactions in this SL-BL system,

### v2

@@ Abstract
The discovery of superconductivity (SC) with critical temperature $T_{c}$ above the boiling point of liquid nitrogen in pressurized La3Ni2O7 has sparked a surge of exploration of high-$T_{c}$ superconductors in the Ruddlesden-Popper (RP) phase nickelates. More recently, the RP phase nickelate La5Ni3O11, which hosts a layered structure with alternating bilayer and single-layer NiO2 planes, has been reported to accommodate SC under pressure, exhibiting a dome-shaped pressure dependence with highest $T_{c}\approx 64$ K, capturing a lot of interest. Here, using density functional theory (DFT) and random phase approximation (RPA) calculations, we systematically study the electronic properties and superconducting mechanism of this material. Our DFT calculations yield a band structure including two nearly decoupled sets of sub-band structures, with one set originating from the bilayer subsystem and the other from the single-layer one. RPA-based analysis demonstrates that SC in this material occurs primarily within the bilayer subsystem exhibiting an $s^{\pm}$ wave pairing symmetry similar to that observed in pressurized La3Ni2O7, while the single-layer subsystem mainly serves as a bridge facilitating the inter-bilayer phase coherence through the interlayer Josephson coupling (IJC). Since the IJC thus attained is extremely weak, it experiences a prominent enhancement under pressure, leading to the increase of the bulk $T_{c}$ with pressure initially. When the pressure is high enough, the $T_{c}$ gradually decreases due to the reduced density of states on the $\gamma$-pocket. In this way, the dome-shaped pressure dependence of $T_{c}$ observed experimentally is naturally understood.
Hybrid nickelate, Pairing mechanism, Interlayer Josephson coupling
@@
The discovery of superconductivity (SC) with critical temperature $T_{c}\approx 80$ K in the Ruddlesden-Popper (RP) phase bilayer nickelate La3Ni2O7 under pressure Sun et al. (2023) has sparked significant experimental and theoretical interest in the field Luo et al. (2023); Huo et al. (2025); Zhang et al. (2023); Lechermann et al. (2023); Luo et al. (2024); Wú et al. (2024); Wang et al. (2024b); Shilenko and Leonov (2023); Yang et al. (2023b); Huang et al. (2023); Christiansson et al. (2023); Shen et al. (2023); Gu et al. (2025); Oh and Zhang (2023); Liu et al. (2023); Liao et al. (2023); Yang et al. (2023a, 2024b); Kaneko et al. (2024); Ouyang et al. (2024); Heier et al. (2024); Zhang et al. (2024d, e); Tian et al. (2024); Ryee et al. (2024); Zhang et al. (2024a); Ni et al. (2025); Lu et al. (2024); Qu et al. (2024); Yang et al. (2024a); Fan et al. (2024b); Sakakibara et al. (2024a); Cao and Yang (2024); Jiang et al. (2024); Chen et al. (2025); Yang et al. (2024c); Zhang et al. (2024c); Hou et al. (2023); Fan et al. (2024a); Wang et al. (2024a); Liu et al. (2024); Zhang et al. (2024b); Jiang et al. (2025); Shao et al. (2024). Subsequently, researchers observed SC in the trilayer nickelate La4Ni3O10 under pressure, and recently detected SC in ambient-pressure (AP) La3Ni2O7 thin films Zhu et al. (2024); Zhang et al. (2025a); Ko et al. (2025); Zhou et al. (2025). These findings further underscore the potential of RP-phase nickelates as high-$T_{c}$ superconductors Chen et al. (2024a); Xu et al. (2025); Sakakibara et al. (2024b); Li et al. (2024c); Hu et al. (2025); Li et al. (2024b); Wang et al. (2024c); Geisler et al. (2024). In RP phase nickelates with the general formula Rn+1NinO3n+1 (R = rare earth), each Ni forms a NiO6 octahedron with corner-sharing O atoms. At AP, bilayer La3Ni2O7 and trilayer La4Ni3O10 manifest octahedral tilting and in-plane lattice anisotropy Sun et al. (2023); Li et al. (2024b); Zhu et al. (2024). Upon increasing pressure, both the octahedral tilting and the in-plane lattice constant disparity gradually diminish and eventually the crystal structures attain higher symmetry ($I4/mmm$ space group) Li et al. (2025).
@@
Furthermore, hybrid RP phase nickelates, formed by alternating stacking different RP phases along the c-axis, also attract widespread investigations, especially the hybridization of the single-layer La2NiO4, the bilayer La3Ni2O7 and the trilayer La4Ni3O10 Zhang et al. (2024f); Chen et al. (2024b); Puphal et al. (2024); Li et al. (2024a); Ouyang et al. (2025). Until now, successful synthesis of 1313 phase (La3Ni2O7)  Chen et al. (2024b); Puphal et al. (2024) and 1212 phase (La5Ni3O11Li et al. (2024a) has been achieved. A previous study reported a superconducting transition onset temperature up to 80 K in alternating stacking single-layer La2NiO4 and trilayer La4Ni3O10, namely 1313 phase Puphal et al. (2024). However, it remains unclear whether the observed high-$T_{c}$ comes from the hybrid 1313 phase or 2222 bilayer phase.
@@
Recently, pressure-induced SC with maximal $T_{c}=$ 64 K was observed in the hybrid 1212 phase, i.e. La5Ni3O11 Shi et al. (2025). Transport and magnetic torque measurements also revealed a density-wave (DW) transition at ambient and low pressures. In contrast to the previously reported DW in La3Ni2O7 and La4Ni3O10, the DW observed here is notably robust against pressure. Intriguingly, SC emerges at around 11.7 GPa and exhibits a dome-shaped pressure-dependent behavior, in contrast to the previously reported right-triangle-like behavior observed in pressurized La3Ni2O7 Li et al. (2025). Previous random-phase approximation (RPA) based study proposed that $d$-wave is the leading pairing symmetry that originates from the La2NiO4 single-layer in the material Zhang et al. (2025b). However, recent studies using combined density functional theory (DFT) and dynamic-mean-field-theory (DMFT) provide another possibility that the La2NiO4 single-layer in La5Ni3O11 is nearly Mott-insulating which does not carry SC Ouyang et al. (2025); LaBollita and Botana (2025), consistent with previous insight that the pure La2NiO4 is an antiferromagnetic (AFM) Mott insulator Yao and Carlson (2007); Freeman et al. (2009). Therefore, a more systematic investigation into the electronic and superconducting properties of the hybrid 1212 phase La5Ni3O11 is urgently needed.
@@
In this paper, we adopt the first-principle DFT calculations to study the electronic structure of La5Ni3O11, followed by a RPA based study to clarify the superconducting mechanism in this material. Based on our DFT band structure, we construct a trilayer $(d_{z^{2}},d_{x^{2}-y^{2}})$-orbital tight-binding (TB) model, in which the hopping integrals between the bilayer subsystem and the single-layer one is very weak, reflecting that the two subsystems are almost decoupled, allowing for the two parts to be treated separately. Our RPA based analysis suggests that the superconducting pairing mainly occurs in the bilayer subsystem, which takes the s±-pattern similar with that in pressurized La3Ni2O7, while the single-layer subsystem mainly serves as a bridge connecting adjacent superconducting bilayer subsystems to establish phase coherence along the c-axis necessary for the bulk SC, through the interlayer Josephson coupling (IJC). As the IJC is extremely weak, it will experience prominent enhancement with the increase of the pressure, leading to enhancement of the $T_{c}$ in the low pressure regime. When the pressure is strong enough, the $T_{c}$ decreases with further enhancement of the pressure due to a reduction in the density of states (DOS) on the $\gamma$-pocket crucial for pairing. Therefore, our study provides a natural understanding of the dome-shaped pressure dependence of the $T_{c}$ observed in the experiment. Furthermore, our findings elucidate the distinct pressure responses of hybrid-phase and pure-phase nickelate superconductors, offering a unified framework that may also be extended to other hybrid superconducting systems.
@@ Results
Figure 1: (color online) Side view of crystal structures of the single-layer (La2NiO4), bilayer (La3Ni2O7) and SL-BL (La5Ni3O11) RP phase nickelates. The blue, grey, and red balls represent lanthanum, nickel, and oxygen atoms, respectively.
La5Ni3O11 is a single-layer-bilayer (SL-BL) stacked nickelate, as shown in Fig. 1. La2NiO4 and La3Ni2O7 stack alternatively in the RP phase, crystallizing into an orthorhombic phase ($Cmmm$ or $Immm$ space group) at AP and transitioning to $P4/mmm$ space group under high pressure (HP) (Shi et al., 2025; Li et al., 2024a). DFT calculated phonon spectra show imaginary frequencies at AP whereas the imaginary frequencies vanish under HP for $P4/mmm$ phase (see Fig.S1 54), demonstrating a structural phase transition under pressure and the structural stability of $P4/mmm$ phase under high pressure. Unlike bulk La3Ni2O7, the out-of-plane Ni-O-Ni angle in La5Ni3O11 remains 180 from AP to HP, without octahedral rotation. Therefore, in La5Ni3O11 each unit cell contains only one Ni atom per layer, in contrast to the cases of bulk La3Ni2O7 and La4Ni3O10 in which each unit cell contains two Ni atoms per layer at AP due to octahedral rotation.
The band structure of the SL-BL stacking La5Ni3O11 at 12 GPa is shown in Fig. 2(a), consistent with previous theoretical results Ouyang et al. (2025); LaBollita and Botana (2025). As in most RP phase nickelates, bands around the Fermi level can be described by the Ni-$e_{g}$ sector. The SL-derived Ni-$d_{z^{2}}$ band lies between the BL-derived bonding and anti-bonding bands. Compared to the band structure at AP (see Fig.S2 54), the BL-derived bonding band is metalized by pressure, leaving the $\gamma$ pocket at the brillouin zone (BZ) corner, as shown in Fig. 2(b). This phenomenon resembles the behavior of pressurized bulk La3Ni2O7.
Figure 2: (color online) Band structure of the DFT and six-orbital TB model for La5Ni3O11 at 12 GPa. (a) DFT band structure and projected DOS of La5Ni3O11. (b) FS in the BZ, with the five pockets labeled. The color scheme in the left half of (b) represents the relative contributions of the $d_{z^{2}}$ and $d_{x^{2}-y^{2}}$ orbital, while the colors scheme on the right half of (b) indicates the relative contributions from Ni atoms in the single-layer and bilayer subsystems. (c) Schematic of La5Ni3O11 lattice of six-orbital TB model. The dashed line denotes the $(\frac{1}{2},\frac{1}{2})$ translation of RP stacking between bilayer and single-layer sublattice. (d) The TB band structure corresponding to (c).
@@
Here $\bm{r}_{i}$ represents the coordinates of site $i$, $\Delta\bm{r}_{x}(\Delta\bm{r}_{y})\in(-2,2)$ represents the hoppings up to the third-nearest neighbor. In this context, $\bm{r}_{i}$ refers to the unit cell index, while a finite $\Delta\bm{r}$ represents the interunit cell separation. The indices $\mu,\nu=1,\cdots,6$ containing the $d_{z^{2}}$ and $d_{x^{2}-y^{2}}$ orbitals of the upper and lower layers of the bilayer subsystem, as well as those of the single-layer subsystem. The SL-BL lattice and corresponding hoppings are depicted in Fig. 2(c), with the hopping integrals $t_{ij}^{\mu\nu}$ provided in Tab. 1. Note that due to the RP stacking in La5Ni3O11, a $(\frac{1}{2},\frac{1}{2})$ translation exists between the bilayer and single-layer sublattice. Fig. 2(d) displays the band structure for TB model, which is in good agreement with DFT results. The obtained FS for this TB model contains five pockets labeled as $\alpha$, $\alpha^{\prime}$, $\beta$, $\gamma$, $\gamma^{\prime}$, as shown in Fig. 2(b). The dominant component orbital of $\gamma$ and $\gamma^{\prime}$ pockets is $d_{z^{2}}$, while the $\alpha$, $\alpha^{\prime}$ and $\beta$ pockets are the mix of $d_{z^{2}}$ and $d_{x^{2}-y^{2}}$ orbitals. From the right half of Fig. 2(b), we can see that the $\alpha^{\prime}$ and $\gamma^{\prime}$ pocket are contributed from single-layer subsystem while the other three pockets from bilayer subsystem.
Table 1: Hopping parameters and site energies of TB model for La5Ni3O11 at 12 GPa in unit of eV. Here, $x$, $z$ and xzxz denotes hopping within and between $d_{x^{2}-y^{2}}$, $d_{z^{2}}$ orbitals, respectively. The abbreviations BI, BO, and SI represent intra-bilayer, inter-bilayer, and single-layer in-plane hoppings, respectively, while BS denotes hoppings between bilayer and single-layer units. The TB model includes hopping processes up to the third-nearest neighbor, i.e., [2,0]. $\epsilon$ denotes the on-site energy.
@@
Figure 3: (color online) TB bands and FS characteristics of the BL and SL subsystems in La5Ni3O11 (SL-BL system) at 12 GPa. (a) Band structure of the SL-BL system (black lines) and that of the isolated BL subsystem (red lines). (b) FS of the isolated BL subsystem, showing that the $\alpha$ pocket is primarily composed of $d_{x^{2}-y^{2}}$ orbitals, the $\beta$ pocket contains contributions from both $d_{x^{2}-y^{2}}$ and $d_{z^{2}}$ orbitals, and the $\gamma$ pocket originates mainly from $d_{z^{2}}$ orbitals. (c) Band structure of the SL-BL system (black lines) and the isolated SL subsystem (red lines). (d) FS of the isolated SL subsystem, showing that the $\alpha^{\prime}$ pocket is primarily composed of $d_{x^{2}-y^{2}}$ orbitals, the $\gamma^{\prime}$ pocket originates mainly from $d_{z^{2}}$ orbitals.
@@
As shown in Tab. 1, the coupling between the single-layer subsystem and the bilayer subsystem is extremely weak, with the strongest hopping integral amplitude between the two parts being $\left|t_{[\frac{1}{2},\frac{1}{2}]}^{z,BS}\right|$=0.0103 eV. The sole effect of such a weak coupling is to equilibrate the chemical potentials of the two subsystems and thereby adjust the electron filling. Consequently, it is reasonable to treat the two subsystems separately.
@@
The decoupling of the two subsystems at the TB Hamiltonian level is achieved by modifying the summation over $\mu$ and $\nu$ in Eq. 1. Specifically, restricting $\mu,\nu\in(1,\cdots,4)$ and electron filling $n\approx 3.1$ which got from SL-BL system yields the Hamiltonian of the bilayer subsystem, while restricting $\mu,\nu\in(5,6)$ and electron filling $n\approx 1.9$ yields that of the single-layer subsystem. A comparison between the band structure of the bilayer subsystem and that of La5Ni3O11 (SL-BL system) is shown in Fig. 3(a), where it can be seen that the bilayer subsystem reproduces the key features of the bilayer-derived bands in the full system. The three FS pockets of the bilayer subsystem, labeled $\alpha$, $\beta$, and $\gamma$, are shown in Fig. 3(b), and are clearly separated from the five pockets observed in Fig. 2(b). Fig. 3(c) shows that the band structure of the single-layer subsystem matches well with the single-layer-derived bands of the full SL-BL system. The two Fermi pockets of the single-layer subsystem, corresponding to the $\alpha^{\prime}$ and $\gamma^{\prime}$ pockets of the full system, are shown in Fig. 3(d).
@@
Figure 4: (color online) (a) Distribution of the largest eigenvalue of the spin susceptibility matrix $\chi^{s}(q)$ in the BZ in the bilayer subsystem for $U=1$ eV and $J_{H}=U/6$. The susceptibility peaks at two inequivalent momenta, denoted as Q1 and Q2, respectively. (b) FS of the bilayer subsystem in the BZ at 12 GPa. As shown in (b), Q1 corresponds to a nesting vector between the $\beta$ and $\gamma$ pockets, while Q2 corresponds to a nesting vector between the $\alpha$ and $\beta$ pockets. (c) The largest pairing eigenvalue $\lambda$ of the various pairing symmetries as function of the interaction strength $U$ with fixed $J_{H}=U/6$. (d) The distributions of the leading $s$-wave pairing gap function on the FS for $U=1.1$ eV, exhibiting an $s^{\pm}$-pattern.
@@
The separation of the single-layer and bilayer subsystems within the TB framework points toward an underlying layer-selective character in this hybrid RP-phase nickelate. At ambient and low pressures, the incorporation of local dynamical correlations via the DMFT modifies the DFT band structure by removing the FS pockets contributed by the single‐layer subsystem, leaving only the three FS pockets contributed by the bilayer subsystem Ouyang et al. (2025); LaBollita and Botana (2025). This observation indicates that the single-layer subsystem is close to a Mott‐insulating state, with possible AFM order present Ouyang et al. (2025), analogous to the bulk La2NiO4 Yao and Carlson (2007); Freeman et al. (2009). At HP, the Mott gap becomes narrower, permitting a small density of mobile carriers that may suppress the AFM order Ouyang et al. (2025). Indeed, the long‐range magnetic order has not been observed experimentally under HP so far Li et al. (2024a). Regardless of whether residual magnetic order persists or not, the single-layer subsystem remains a “bad metal” and can hardly carry SC Ouyang et al. (2025). This suggests that superconductivity is more likely to originate from the bilayer subsystem. Since DFT+DMFT indicates that electronic correlation in the bilayer subsystem is comparatively weaker, we will focus our subsequent RPA analysis on the bilayer subsystem to study its pairing instability.
@@
Given that superconductivity is likely to arise predominantly from the bilayer subsystem, we adopt the following multi-orbital Hubbard interaction to investigate the SC driven by electron interactions in this bilayer subsystem,